1. Let X.. xs ^ N(-1,4) and Y.. y, N(0.1) be independent. Using properties of the normal distribution, derive the distribution of the following random variables (b) Wa = 2.1(X + 1)
Answer #1

(a)

We know that the linear combination of Normal random variable is a Normal random variable.

Thus, W1 is a Normal random variable.

E(W1) = E(2X1 + 4Y2 - X2) = 2E(X1) + 4E(Y2) - E(X2) = 2 * - 1 + 4 * 0 - (-1) = -1

Var(X1) = Var(2X1 + 4Y2 - X2) = 22Var(X1) + 42Var(Y2) + (-1)2 Var(X2) = 4 * 4 + 16 * 1 + 1 * 4 = 36

W1 ~ N(-1, 36)

(b)

We know that the linear combination of Normal random variable is a Normal random variable.

Thus, W2 is a Normal random variable.

3 2 3 2 3 2 2

=> E[W2] = 0

3 3 3 2 2 Var (15) = Var VarXi

W2 ~ N(0, 3)

(c)

Let us consider the variable Z1 = (X1 + 1)/2

E[(X1 + 1)/2] = (E(X1) + 1 )/2 = (-1 + 1) / 2 = 0

Var[(X1 + 1)/2] = (Var(X1) )/22 = 4 / 4 = 1

Thus, Z1 = (X1 + 1)/2 ~ N(0, 1)

and Y3 ~ N(0,1)

We know that sum of squares of k standard normal variables follow Chi square distribution with k degree of freedom.

Thus, W3 = Z12 + Y32 follows Chi square distribution with 2 degree of freedom.

(d)

We know that sum of squares of k standard normal variables follow Chi square distribution with k degree of freedom.

sum_{i=1}^{4} Y_i^2 follow Chi square distribution with 4 degree of freedom.

Student's t-distribution with ν degrees of freedom can be defined as the distribution of the random variable T with

T = Z/ sqrt{V/v}

where

  • Z is a standard normal with expected value 0 and variance 1;
  • V has a chi-squared distribution with ν degrees of freedom;
  • Z and V are independent.

Now,

W_4 = rac{2Y_5}{sqrt{sum_{i=1}^{4} Y_i^2}} = rac{Y_5}{sqrt{sum_{i=1}^{4} Y_i^2}/4}

As, Y5 ~ N(0,1) and sum_{i=1}^{4} Y_i^2 follow Chi square distribution with 4 degree of freedom.

W4 follows Student's t-distribution with 4 degrees of freedom